OpenStudy (anonymous):

In square ABCD, MC = 12 ft. Which is AB?

OpenStudy (sriram):

what is M??

OpenStudy (sriram):

2*12/(2)^1/2 =12*(2)^1/2

OpenStudy (anonymous):

\[AC^2+BC^2=AC^2\]Ac=24 \[24^2=AB^2+BC^2\]AB=BC \[2AB^2=576\]\[AB=\sqrt{\frac{576}{2}}\] AB=16.98 if you like my answer, click good answer,TY

OpenStudy (anonymous):

AC=12*2 AC=24

OpenStudy (sriram):

nancy lam i m jst honoured to give u ur 300th medal bt ur method is quite lenthy use trigo

OpenStudy (anonymous):

\[12\sqrt{2}\]